# Localize, Understand, Collaborate: Semantic-Aware Dragging via Intention Reasoner

### v1

@@ Abstract
Flexible and accurate drag-based editing is a challenging task that has recently garnered significant attention. Current methods typically model this problem as automatically learning "how to drag" through point dragging and often produce one deterministic estimation, which presents two key limitations: 1) Overlooking the inherently ill-posed nature of drag-based editing, where multiple results may correspond to a given input, as illustrated in Fig. 1; 2) Ignoring the constraint of image quality, which may lead to unexpected distortion. To alleviate this, we propose LucidDrag, which shifts the focus from "how to drag" to a paradigm of "what-then-how." LucidDrag comprises an intention reasoner and a collaborative guidance sampling mechanism. The former infers several optimal editing strategies, identifying what content and what semantic direction to be edited. Based on the former, the latter addresses "how to drag" by collaboratively integrating existing editing guidance with the newly proposed semantic guidance and quality guidance. Specifically, semantic guidance is derived by establishing a semantic editing direction based on reasoned intentions, while quality guidance is achieved through classifier guidance using an image fidelity discriminator. Both qualitative and quantitative comparisons demonstrate the superiority of LucidDrag over previous methods. The code will be released.
@@ 1 Introduction
To address this, recent studies  [15, 45, 48] define a new task called drag-based editing, which incorporates dragging points as conditions. These studies specifically regard drag-based editing as the problem of "how to drag" and tackle it by designing an editing guidance loss, enabling the model to implicitly learn the appropriate solutions. Despite their considerable success [45, 40, 41], these methods have two limitations: Firstly, they neglect the inherent ambiguity of semantic intention. Drag-based editing is an ill-posed problem. As depicted in Fig. 1, the drag points starting from the horse’s head and ending at its upper right can indicate various semantic intentions, such as "make the neck longer," "raise the head," or "bring the two horses closer." However, existing methods mainly follow point dragging principles [48, 40], focusing on positional movement by constraining feature correlation between the source and target points. This current position optimization strategy inherently overlooks semantic diversity, making it challenging to generate images with precise semantic perception. Secondly, they overlook the preservation of the overall image quality. Current methods prioritize editing accuracy, often neglecting the overall image quality. Some methods [40, 41] utilize score-based classifier guidance for image editing, which can cause mismatches between the distribution of the edited image and the input image, compromising image fidelity.
@@
In this study, we divide the drag-based editing task into two steps. We introduce a preliminary step, "what to drag", to determine the specific content and semantic direction for editing before addressing "how to drag". That is, we shift the focus from "how to drag" to a paradigm of "what-then-how". As shown in the first row of images in Fig. 1, before editing, we need to determine what we are going to edit, such as the horse’s head, and the semantic strategy for editing it towards the upper right. For example, we could make the horse lift its head, elongate its neck, or shorten the distance between the two horses. With the "what to drag" information established, we can then proceed to address "how to drag". To achieve this, we construct an intention reasoner that integrates a Large Language-Vision Model(LVLM) and a Large Language Model(LLM) to deduce possible intentions. As illustrated in Fig. 1, given an input image and drag points that begin at the horse’s face and end in its upper-right, the intention reasoner acts as an AI agent to infer potential intentions, subsequently providing corresponding source and target prompts. Once the intention is determined, we then address “how to drag” by injecting the reasoned semantic information into the model by developing collaborative guidance sampling, which integrates editing guidance with the proposed semantic guidance and quality guidance. Specifically, the semantic guidance is derived from the source and target prompts determined by the intention reasoner. The asymmetric prompts establish a clear semantic editing direction towards the target intention. Additionally, a discriminator is employed as the score function to provide quality guidance. The quality gradient is generated based on image fidelity and incorporated into the model via the classifier guidance mechanism. For the editing guidance, we follow previous work [40] to maximize the feature correspondence between the source and target positions.
@@ 2.3 Visual Programming
As LLMs [44, 7, 52] and LVLMs [66, 12, 59] demonstrated remarkable emergency abilities [1, 55], researches [19, 34, 53] explore to leverage them for planning and reasoning in multi-modal image generation. For example, ChatEdit [10] utilizes pre-trained language models to track the user intents. Some approaches [19, 34, 53] augment the original prompt through paraphrasing. Recent, visual programmer methods [23, 24, 8] translate complex input prompts into programmatic operations and data flows. Despite the effectiveness of these strategies in augmenting the input text instructions, they overlook the capacity to reason within visual-modal instructions, such as the dragging points in drag editing tasks. In contrast, our LucidDrag addresses this gap by integrating LVLM and LLM.
@@ 2.4 Image Editing
Image editing aims to manipulate an image according to specific conditions. Prior methodologies have primarily relied on text conditions. For example, [10, 2] manipulate images in the StyleGAN [29] latent space by learning an edit direction. Motivated by the success of the diffusion model [27], state-of-the-art methods extend their exploration into diffusion-based image editing by exploring the initial noise [11, 65, 38], attention maps [6, 26, 56, 16], or prompts [33, 39, 14, 4]. Recently, DragGAN [45] explores a novel editing scheme that drags any points of the image to reach target points with the help of StyleGAN [29] latent space. FreeDrag [35] improves point tracking by introducing adaptive feature updating and backtracking. Readout Guidance [37] solves the challenging task by leveraging the video dataset. The following works  [40, 41] utilize the feature correspondence to direct the editing process. Nevertheless, they only use the drag points as a control, which is insufficient due to the potential diversity in semantic intentions. In addition, existing drag edit methods primarily focus on designing algorithms to manipulate content, overlooking the potential distortions, which may reduce the overall image quality. In contrast, our approach introduces an intention reasoner to achieve semantic-aware editing. Moreover, we incorporate quality guidance to avoid image distortions by constraining image quality explicitly.
@@ 3 Method
In this section, we introduce LucidDrag, a unified framework for Drag Manipulation via Localing, Understanding, Collaborate Guiding. Within our framework, image drag editing is decomposed into two stages. Firstly, the Intention Reasoner translates the user-drag points into potential semantic intentions and generates corresponding source prompt and target prompt, thus solving the problem of “what to edit”. Then, the Collaborative Guidance Sampling is designed to facilitate image editing. Concretely, prompts generated from the intention reasoner are utilized in the DDIM Inversion and Diffusion U-Net, providing semantic guidance for the generation of semantically controllable results. Besides, LucidDrag employs a discriminator to provide quality guidance for improved image fidelity. The semantic guidance and the quality guidance collaborate with the original editing guidance, offering a novel perspective of “how to edit”. We elaborate on more details of our techniques below.
@@ 3.1 Intention Reasoner
As shown in Fig. 2, each point condition may encompass various semantic intentions. For instance, it may represent non-rigid manipulation, such as “looking down”, while it may also represent rigid manipulation, such as “moving to the left”. Our LLM-driven reasoner is designed to discern potential semantic intentions to facilitate semantic-aware drag-based editing. We leverage the capabilities of the large language model, GPT 3.5 [44], acting as an AI agent to reason the possible intentions. We take the generated description of the object of interest $O$, the original image caption $C$, and drag points $P$ as input. Then, we prompt the LLM to generate $N$ possible intentions, i.e. $D=LLM(O,C,P)$, where each output sample can have different intentions or levels of complexity. Specifically, $D=\left\{\left(d_{j},P\left(d_{j}\right)\right)\right\}_{j=1}^{N}$, where $d_{j}=\left\{i_{j},p_{j}^{s},p_{j}^{t}\right\}$ is the generated text output. $i_{j}$, $p_{j}^{s}$, $p_{j}^{t}$ represents the predicted intention, the predicted description of the source image (source prompt), and the predicted description of the target image (target prompt), respectively. $P\left(d_{j}\right)$ is the corresponding confidence probabilities of the $j\text{-}th$ generate text output. Finally, we can select $n$ outputs $\left\{d_{s^{*}}\right\}_{s=1}^{n}$ by sampling based on the confidence probabilities:
@@ 3.2 Collaborative Guidance Sampling
As shown in Fig. 2, the objective of collaborative guidance sampling is to modify the intended content while ensuring the preservation of irrelevant components. The input image is inverted to noise $z_{T}^{gud}$ through DDIM Inversion [50, 39]. During the inversion process, the intermediate noise $z_{t}^{gud}$, along with the corresponding key $k_{t}^{gud}$, and value $v_{t}^{gud}$ of the self-attention layer, are recorded in the memory bank, which serves in guiding subsequent generation process. Subsequently, we generate the edited images employing collaborative guidance sampling which incorporates three fundamental components: semantic guidance, quality guidance, and editing guidance. Each of these components contributes to the overall editing process distinctively, thereby ensuring a balanced and comprehensive approach to image editing.
@@ 4.2 Comparisons
Figure 4: Qualitative comparison between our LucidDrag and other methods in drag-based editing.
@@
We also present qualitative results in Fig. 4. DragDiffusion and FreeDrag has difficulty in accurately dragging corresponding contents to designated target locations. Although DragonDiffusion and DiffEditor can better recognize handle points and achieve more precise point movement, they tend to introduce artifacts, thereby reducing image quality. In contrast, LucidDrag demonstrates superior drag control by understanding potential intentions and providing semantic guidance. Furthermore, LucidDrag generates images with greater fidelity due to the explicit quality guidance it provides.
@@ 5 Conclusion
In this work, we identify the limitations of previous drag editing models in understanding semantic intentions and generating high-quality edited images. In response, we design a novel framework called LucidDrag, which involves an intention reasoner to clarify possible intentions and a collaborative guidance sampling mechanism that incorporates explicit semantic guidance and quality guidance. LuciDrag has merits in: i) adequate understanding of semantic intention, improving semantic perception ability and diversity of the generated images; ii) enhanced dragging performance, including improved drag accuracy and image quality. Extensive results show the efficiency of our approach and the potential for further advancements in semantic-aware drag-based editing.
@@ Appendix A Appendix
  • §A.4 More Comparisons.
  • §A.5 Visualizations of the Guidance.
  • §A.6 Additional Ablation Study for Hyperparameters.
@@ A.1 Additional Preliminaries of Diffusion Model
where $\varnothing=\psi($“”)))) is the embedding of a null text. $\varepsilon_{\theta}(z_{t},t,y)$ represents the conditional predictions. $w$ is the guidance scale parameter.
@@ A.4 More Comparisons
DiffEditor $24.41\pm 2.46$ $27.52\pm 4.32$ $34.94\pm 4.63$ $38.42\pm 6.45$ $25.62\pm 2.14$ $24.60\pm 2.43$ $24.77\pm 2.13$ $6.81\pm 0.44$ $17.71\pm 3.07$ $35.32\pm 2.55$ $26.01\pm 0.81$
Ours $\textbf{19.71}\pm 0.31$ $\textbf{17.76}\pm 1.66$ $\textbf{26.74}\pm 1.08$ $\textbf{26.67}\pm 1.87$ $\textbf{22.74}\pm 2.45$ $\textbf{23.87}\pm 0.91$ $\textbf{20.70}\pm 0.86$ $\textbf{5.56}\pm 0.98$ $\textbf{14.34}\pm 0.28$ $\textbf{26.54}\pm 1.69$ $\textbf{20.46}\pm 0.77$
@@
Figure 7: More qualitative comparison on conten dragging.
@@
More Qualitative Results
To begin, we present additional qualitative results that demonstrate the semantic awareness capability of our proposed method. As illustrated in Fig. 8, our approach is capable of generating a diverse range of output images that are highly faithful to the intentions that have been deduced.
Subsequently, further comparison results for the content dragging task are presented in Fig. 7. These results, in conjunction with those presented in Fig. 4 of the main paper, substantiate the superiority of our method in terms of both drag-based editing performance and image quality maintenance. The extensive qualitative experiments serve to further substantiate the generality of our approach.
A.5 Visualizations of the Guidance
@@
Figure 9: Visualization of the quality guidance an editing guidance.
A.6 Additional Ablation Study for Hyperparameters
@@
A.7 Social Impacts

### v2

@@ Abstract
Flexible and accurate drag-based editing is a challenging task that has recently garnered significant attention. Current methods typically model this problem as automatically learning "how to drag" through point dragging and often produce one deterministic estimation, which presents two key limitations: 1) Overlooking the inherently ill-posed nature of drag-based editing, where multiple results may correspond to a given input, as illustrated in Fig. 1; 2) Ignoring the constraint of image quality, which may lead to unexpected distortion. To alleviate this, we propose LucidDrag, which shifts the focus from "how to drag" to "what-then-how" paradigm. LucidDrag comprises an intention reasoner and a collaborative guidance sampling mechanism. The former infers several optimal editing strategies, identifying what content and what semantic direction to be edited. Based on the former, the latter addresses "how to drag" by collaboratively integrating existing editing guidance with the newly proposed semantic guidance and quality guidance. Specifically, semantic guidance is derived by establishing a semantic editing direction based on reasoned intentions, while quality guidance is achieved through classifier guidance using an image fidelity discriminator. Both qualitative and quantitative comparisons demonstrate the superiority of LucidDrag over previous methods. Code is available at: https://github.com/cuixing100876/LucidDrag-NeurIPS2024.
@@ 1 Introduction
To address this, recent studies  [14, 50, 53] define a new task called drag-based editing, which incorporates dragging points as conditions. These studies specifically regard drag-based editing as the problem of "how to drag" and tackle it by designing an editing guidance loss, enabling the model to implicitly learn the appropriate solutions. Despite their considerable success [50, 45, 46], these methods have two limitations: Firstly, they neglect the inherent ambiguity of semantic intention. Drag-based editing is an ill-posed problem. As depicted in Fig. 1, the drag points starting from the horse’s head and ending at its upper right can indicate various semantic intentions, such as "make the neck longer," "raise the head," or "bring the two horses closer." However, existing methods mainly follow point dragging principles [53, 45], focusing on positional movement by constraining feature correlation between the source and target points. This current position optimization strategy inherently overlooks semantic diversity, making it challenging to generate images with precise semantic perception. Secondly, they overlook the preservation of the overall image quality. Current methods prioritize editing accuracy while neglects the overall image quality. Some methods [45, 46] utilize score-based classifier guidance for image editing, which can cause mismatches between the distribution of the edited image and the input image, compromising image fidelity.
@@
In this study, we divide the drag-based editing task into two steps. We introduce a preliminary step, "what to drag", to determine the specific content and semantic direction for editing before addressing "how to drag". That is, we shift the focus from "how to drag" to a paradigm of "what-then-how". As shown in the first row of images in Fig. 1, before editing, we need to determine what we are going to edit, such as the horse’s head, and the semantic strategy for editing it towards the upper right. For example, we could make the horse lift its head, elongate its neck, or shorten the distance between the two horses. With the "what to drag" information established, we can then proceed to address "how to drag". To achieve this, we construct an intention reasoner that integrates a Large Language-Vision Model (LVLM) and a Large Language Model (LLM) to deduce possible intentions. As illustrated in Fig. 1, given an input image and drag points that begin at the horse’s face and end in its upper-right, the intention reasoner acts as an AI agent to infer potential intentions, subsequently providing corresponding source and target prompts. Once the intention is determined, we then address "how to drag" by injecting the reasoned semantic information into the model by developing collaborative guidance sampling, which integrates editing guidance with the proposed semantic guidance and quality guidance. Specifically, the semantic guidance is derived from the source and target prompts determined by the intention reasoner. The asymmetric prompts establish a clear semantic editing direction towards the target intention. Additionally, a discriminator is employed as the score function to provide quality guidance. The quality gradient is generated based on image fidelity and incorporated into the model via the classifier guidance mechanism. For the editing guidance, we follow previous work [45] to maximize the feature correspondence between the source and target positions.
@@ 2.3 Visual Programming
As LLMs [49, 7, 58] and LVLMs [74, 11, 67] demonstrated remarkable emergency abilities [1, 62, 40, 41], researches [18, 37, 59] explore to leverage them for planning and reasoning in multi-modal image generation. For example, ChatEdit [9] utilizes pre-trained language models to track the user intents. Some approaches [18, 37, 59] augment the original prompt through paraphrasing. Recent, visual programmer methods [22, 24] translate complex input prompts into programmatic operations and data flows. Despite the effectiveness of these strategies in augmenting the input text instructions, they overlook the capacity to reason within visual-modal instructions, such as the dragging points in drag editing tasks. In contrast, our LucidDrag addresses this gap by integrating LVLM and LLM.
@@ 2.4 Image Editing
Image editing aims to manipulate an image according to specific conditions. Prior methodologies [33, 36, 34] can only manipulate specific attributes. The prevailing approaches have primarily relied on text conditions. For example, [9, 2] manipulate images in the GAN [29] latent space by learning an edit direction. Motivated by the success of the diffusion model [27, 60], state-of-the-art methods extend their exploration into diffusion-based image editing by exploring the initial noise [10, 73, 43], attention maps [6, 26, 63, 15], or prompts [35, 44, 13, 4, 57]. Recently, DragGAN [50] explores a novel editing scheme that drags any points of the image to reach target points with the help of StyleGAN [29] latent space. FreeDrag [38] improves point tracking by introducing adaptive feature updating and backtracking. Readout Guidance [42] solves the challenging task by leveraging the video dataset. The following works  [45, 46] utilize the feature correspondence to direct the editing process. Nevertheless, they only use the drag points as a control, which is insufficient due to the potential diversity in semantic intentions. In contrast, our approach introduces an intention reasoner to achieve semantic-aware editing. The intention reasoner can reduce cognitive load, handle vague requests, and discover potential needs. Additionally, it generates precise descriptions automatically, ensuring accurate and consistent manipulations.
@@ 3 Method
In this section, we introduce LucidDrag, a unified framework for Drag Manipulation via Localing, Understanding, Collaborate Guiding. Within our framework, image drag editing is decomposed into two stages. Firstly, the Intention Reasoner translates the user-drag points into potential semantic intentions and generates corresponding source prompt and target prompt, thus solving the problem of "what to edit". Then, the Collaborative Guidance Sampling is designed to facilitate image editing. Concretely, prompts generated from the intention reasoner are utilized in the DDIM Inversion and Diffusion U-Net, providing semantic guidance for the generation of semantically controllable results. Besides, LucidDrag employs a discriminator to provide quality guidance for improved image fidelity. The semantic guidance and the quality guidance collaborate with the original editing guidance, offering a novel perspective of "how to edit". We elaborate on more details of our techniques below.
@@ 3.1 Intention Reasoner
As shown in Fig. 2, each point condition may encompass various semantic intentions. For instance, it may represent non-rigid manipulation, such as "looking down", while it may also represent rigid manipulation, such as "moving to the left". Our LLM-driven reasoner is designed to discern potential semantic intentions to facilitate semantic-aware drag-based editing. We leverage the capabilities of the large language model, GPT 3.5 [49], acting as an AI agent to reason the possible intentions. We take the generated description of the object of interest $O$, the original image caption $C$, and drag points $P$ as input. Then, we prompt the LLM with in-context examples to generate $N$ possible intentions, i.e. $D=LLM(O,C,P)$, where each output sample can have different intentions or levels of complexity. Specifically, $D=\left\{\left(d_{j},P\left(d_{j}\right)\right)\right\}_{j=1}^{N}$, where $d_{j}=\left\{i_{j},p_{j}^{s},p_{j}^{t}\right\}$ is the generated text output. $i_{j}$, $p_{j}^{s}$, $p_{j}^{t}$ represents the predicted intention, the predicted description of the source image (source prompt), and the predicted description of the target image (target prompt), respectively. $P\left(d_{j}\right)$ is the corresponding confidence probabilities of the $j\text{-}th$ generate text output. Finally, we can select $n$ outputs $\left\{d_{s^{*}}\right\}_{s=1}^{n}$ by sampling based on the confidence probabilities. The confidence probability reflects the quality of the output. A higher confidence probability indicates that the intention is more reasonable, leading to better editing results.
@@ 3.2 Collaborative Guidance Sampling
As shown in Fig. 2, the objective of collaborative guidance sampling is to modify the intended content while ensuring the preservation of irrelevant components. The input image is inverted to noise $z_{T}^{gen}$ through DDIM Inversion [55, 44]. During the inversion process, the intermediate noise $z_{t}^{gud}$, along with the corresponding key $k_{t}^{gud}$, and value $v_{t}^{gud}$ of the self-attention layer, are recorded in the memory bank, which serves in guiding subsequent generation process. Subsequently, we generate the edited images employing collaborative guidance sampling which incorporates three fundamental components: semantic guidance, quality guidance, and editing guidance. Each of these components contributes to the overall editing process distinctively, thereby ensuring a balanced and comprehensive approach to image editing.
@@ 4.2 Comparisons
Qualitative comparison between our LucidDrag and other methods in drag-based editing.
@@
We also present qualitative results in Fig. 4.2. DragDiffusion and FreeDrag has difficulty in accurately dragging corresponding contents to designated target locations. Although DragonDiffusion and DiffEditor can better recognize handle points and achieve more precise point movement, they tend to introduce artifacts, thereby reducing image quality. In contrast, LucidDrag demonstrates superior drag control by understanding potential intentions and providing semantic guidance. Furthermore, LucidDrag generates images with greater fidelity due to the explicit quality guidance it provides.
@@ 5 Conclusion
In this work, we identify the limitations of previous drag editing models in understanding semantic intentions and generating high-quality edited images. In response, we design a novel framework called LucidDrag, which involves an intention reasoner to clarify possible intentions and a collaborative guidance sampling mechanism that incorporates explicit semantic guidance and quality guidance. LucidDrag excels in: i) adequate understanding of semantic intention, improving semantic perception ability and diversity of the generated images; ii) enhanced dragging performance, including improved drag accuracy and image quality. Extensive results show the efficiency of our approach and the potential for further advancements in semantic-aware drag-based editing.
6 Limitations
Although our method is capable of achieving semantic-aware drag-based editing without the need for training, there are some limitations. i) Complex objects are challenging to drag, and unexpected deformation can occur over long distances. This may be attributed to difficulties in comprehending the intricate nature of the object or inaccurate object tracking. In future work, we will investigate the potential for further improvements in performance by utilizing more powerful image generation models and incorporating comprehensive intention understanding. ii) Hyperparameters involved in the editing process will affect the editing results. In future work, we intend to utilize LLM as an agent to automatically determine model hyperparameters to enhance the editing performance.
Acknowledgement This research is sponsored by National Natural Science Foundation of China (Grant No. 62306041), Beijing Nova Program (Grant No. Z211100002121106, 20230484488), CCF-DiDi GAIA Collaborative Research Funds (Grant No. 202414), and Beijing Municipal Science & Technology Commission (Z231100007423015).
@@ Appendix A Appendix
  • §A.4 More Comparsions.
  • §A.5 More Analysis.
@@ A.1 Additional Preliminaries of Diffusion Model
where $\varnothing=\psi($"")))) is the embedding of a null text. $\varepsilon_{\theta}(z_{t},t,y)$ represents the conditional predictions. $w$ is the guidance scale parameter.
@@ A.4 More Comparisons
DiffEditor $24.41$ $27.52$ $34.94$ $38.42$ $25.62$ $24.60$ $24.77$ $6.81$ $17.71$ $35.32$ $26.01$
Ours 19.71 17.76 26.74 26.67 22.74 23.87 20.70 5.56 14.34 26.54 20.46
@@
Figure 6: More qualitative comparison on content dragging.
More Qualitative Results
More comparison results for the content dragging task are presented in Fig. 6. These results, in conjunction with those presented in Fig. 4.2 of the main paper, substantiate the superiority of our method in terms of both drag-based editing performance and image quality maintenance, and also demonstrate the generality of our approach.
Subsequently, we present additional qualitative results that demonstrate the semantic awareness capability of our proposed method. As illustrated in Fig. 7, our approach is capable of generating a diverse range of output images that are highly faithful to the intentions that have been deduced.
A.5 More Analysis
Analysis of different LVLMs and LLMs in Intention Reasoner
We conduct experiments to examine the performance of different LVLMs and LLMs in the Intention Reasoner module. Specifically, we utilize Osprey [67] and Ferret [66] for LVLM and Vicuna [7], LLama3 [58], and GPT 3.5 [49] for LLM. We test various combinations, with Osprey+GPT3.5 being the default setting in our paper. As shown in Table 6, all combinations outperform the experiment without the Intention Reasoner, confirming its reliability. This reliability stems two-fold: the LVLMs are trained with large-scale point-level labeled data and can easily achieve point-level understanding [67]. Therefore, they can understand the user-given points without further fine-tuning. For the LLMs, state-of-the-art LLMs have been proven to possess strong spatial reasoning abilities [23], enabling them to deduce reasonable intentions.
@@ A.5 More Analysis
Table 6: Results with different LVLMs and LLMs
w/o Intention Reasoner Ferret+ Vicuna Ferret+ LLama3 Ferret+ GPT3.5 Osprey+ Vicuna Osprey+ LLama3 Osprey+ GPT3.5 (Ours)
Mean Distance ($\downarrow$) 23.66 22.49 21.96 20.65 20.84 20.48 20.46
GScore ($\uparrow$) 6.76 7.12 7.11 7.35 7.27 7.13 7.37
Analysis of the confidence probabilities
We analyze the confidence probabilities of the intention reasoner. As shown in Fig. 8, the confidence probability reflects the quality of the output text in LLM. A higher confidence probability indicates that the intention of the output is more reasonable, leading to better editing results.
Figure 8: Analysis of confidence probabilities. A higher confidence probability indicates that the intention of the output is more reasonable, leading to better editing results.
Analysis of the guidance
@@
Figure 9: Visualization of the quality guidance and editing guidance.
@@
The efficiency of different methods
We present the efficiency of different methods in Table 7. Our method has a relatively small inference time and comparable memory requirements.
Table 7: Efficiency of different methods.
DragDiffusion FreeDrag DragonDiffusion DiffEditor Ours
Time (s) $\downarrow$ 80 92 30 35 48
Memory (GB) $\downarrow$ 12.8 13.1 15.7 15.7 15.8
Bad case
As illustrated in Fig. 11, some complex objects are challenging to drag and some unexpected deformation occurs when dragging over long distances. This may be attributed to difficulties in comprehending the intricate nature of the object or inaccurate object tracking.
Figure 11: Bad case of our LucidDrag.
A.6 Social Impacts